OpenStudy (anonymous):

Ramona deposited $4,190.51 into a savings account with an interest rate of 5.2% compounded twice a year. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

9000= 4,190.51 * 1.052^n where n = 6 months ln 9000 = ln (4,190.51 * 1.052^n) ln 9000 = ln 4,190.51 + ln 1.052^n ln 9000 - ln 4,190.51 = n ln 1.052 (ln 9000 - ln 4,190.51) / ln 1.052) = n n is about 15.08. but n is 6 months so to convert to years multiply by 2 2n = 30.16 years so it will take 30.16 years compounded twice a year at 5.2% interest.

OpenStudy (barrycarter):

I'm pretty sure 5.2% compounded twice a year means 5.2% for the whole year, or 2.6% every 6 months.

OpenStudy (anonymous):

it should be 15.08/2 to go from months to years, I messed up. whatever it is, just use ln to figure out the problem.

OpenStudy (anonymous):

What is the answer?